# Tony Lindeberg

### v1

@@ 1 Introduction
A conceptual question that one could then ask is if the ability of perceptual system to derive seemingly stable representations of external objects in the 3-D environment would lead to constraints regarding the variabilities in the shapes of the receptive fields at the earliest levels in the visual processing hierarchy. In a theoretically motivated and axiomatically determined normative theory of visual receptive fields (Lindeberg 2013a; 2021), idealized models for the receptive fields of simple cells in the primary visual cortex have specifically been derived, that lead to provable covariance properties under spatial affine transformations, and regarding extensions to time-dependent spatio-temporal image data also provable covariance properties under Galilean transformations, to handle the variabilities in the relative motion between the objects in the world and the observer, see Lindeberg (2023; 2024c) for more in-depth treatments regarding the importance of covariance properties for visual receptive fields under geometric image transformations.
@@ 3.1.1 Monocular locally linearized perspective projection
3.1.2 Binocular locally linearized perspective projection
@@ 6.2 Possible extensions to a 4-D variability to enable matching of affine Gaussian derivative responses
are computed, to instead be allowed to be computed in directions in the spatial domain that do not necessarily have to be aligned with either of the eigenvectors of the spatial covariance matrix $\Sigma$. Then, such extended models for the spatial receptive fields, which will span a four-dimensional variability, will allow for expanded covariance properties over the full group of non-singular spatial affine transformations, according to the treatment in (Lindeberg 2024c Section 3.5).
@@ A.1.6 Relationships for the angular parameters $\alpha$, $\beta$, $\varphi$ and $\psi$
$\begin{split}\psi&=\arctan\left(\frac{C}{S}\right)+n_{\psi}\,\pi,\end{split}$ (97)

### v2

@@ 1 Introduction
A conceptual question that one could then ask is if the ability of perceptual system to derive seemingly stable representations of external objects in the 3-D environment would lead to constraints regarding the variabilities in the shapes of the receptive fields at the earliest levels in the visual processing hierarchy. In a theoretically motivated and axiomatically determined normative theory of visual receptive fields (Lindeberg 2013a; 2021), idealized models for the receptive fields of simple cells in the primary visual cortex (V1) have specifically been derived, that lead to provable covariance properties under spatial affine transformations, and regarding extensions to time-dependent spatio-temporal image data also provable covariance properties under Galilean transformations, to handle the variabilities in the relative motion between the objects in the world and the observer, see Lindeberg (2023; 2024c) for more in-depth treatments regarding the importance of covariance properties for visual receptive fields under geometric image transformations.
@@
The presented results are aimed at providing a theoretical understanding of basic covariance properties for the earliest layers of receptive fields in computational vision, as well as for generating hypotheses about the computational functions in the earliest levels of visual processing in biological vision.
@@ 3.1.1 Monocular locally linearized perspective projection
3.1.2 Binocular locally linearized projective projection
@@ 5 Degrees of freedom in the affine Gaussian derivative model for visual receptive fields
Figure 8: Variability in the orientation of directional derivatives of affine Gaussian derivative receptive fields (for $\sigma_{1}=4$ and $\sigma_{2}=1$), with the orientation angle $\varphi$ increasing from left to right. (top row) First-order directional derivatives of affine Gaussian kernels. (bottom row) Second-order directional derivatives of affine Gaussian kernels. (Horizontal axes: image coordinate $x_{1}\in[-10,10]$. Vertical axes: image coordinate $x_{2}\in[-10,10]$.)
@@ 6.2 Possible extensions to a 4-D variability to enable matching of affine Gaussian derivative responses
An operational explanation why affine covariance cannot be achieved for the regular affine Gaussian derivative kernels, with the orientation for computing the directional derivative of that regular affine Gaussian derivative kernel being required to be aligned with a principal direction of the covariance matrix of the affine Gaussian kernel, is that if the affine image transformation comprises a non-uniform scaling transformation in a direction that not coincides with the principal directions of the spatial covariance matrix in the affine Gaussian smoothing kernel, then the transformed receptive field of a regular affine Gaussian derivative kernel, will be no longer within the definition of such a regular affine Gaussian directional derivative kernel.
Instead, to enable full affine covariance, the image orientation of the directional derivative operator in the transformed affine Gaussian derivative kernel must be allowed to be different from the restriction of the image direction for the directional derivative being aligned to either of the principal directions of the affine Gaussian smoothing kernel.
@@
are computed, to instead be allowed to be computed in directions in the spatial domain that do not necessarily have to be aligned with either of the eigenvectors of the spatial covariance matrix $\Sigma$, see Figure 8 for examples of such receptive fields for fixed values of the spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$. Then, such extended models for the spatial receptive fields, which will span a four-dimensional variability, will allow for expanded covariance properties over the full group of non-singular spatial affine transformations, according to the treatment in (Lindeberg 2024c Section 3.5).
@@ A.1.6 Relationships for the angular parameters $\alpha$, $\beta$, $\varphi$ and $\psi$
$\begin{split}\psi&=\arctan\left(\frac{S}{C}\right)+n_{\psi}\,\pi,\end{split}$ (97)